# Touch in the Wild: Learning Fine-Grained Manipulation with a Portable Visuo-Tactile Gripper

### v1

@@ 5.1 Large-Scale Visuo-Tactile Data and Pretraining
We evaluate the quality of the learned representations from our pretrained visuo-tactile encoder through two complementary qualitative analyses. First, we assess cross-modal reconstruction by providing the encoder with partially masked tactile and RGB images, then measuring its ability to reconstruct the missing tactile inputs. This evaluates whether the model has learned meaningful visuo-tactile associations and can infer tactile information from visual cues, both in in-distribution (seen during training) and out-of-distribution (unseen during training) environments. Second, we examine the encoder’s attention mechanisms by visualizing self-attention maps from the final layer of the Vision Transformer (ViT). This analysis reveals whether the model consistently attends to relevant contact regions in RGB images, even across varied and unstructured scenarios. Fig. 5 illustrates qualitative results from four representative tasks: two from in-distribution and two from out-of-distribution test data.
Figure 5: Qualitative Results of Pretraining. We present four examples showcasing the results of our visuo-tactile pretrained encoder, highlighting both tactile image reconstruction and ViT self-attention heatmaps. For tactile image reconstruction, the encoder successfully reconstructs tactile images for both in-distribution and out-of-distribution inputs. Additionally, we observe that the vision encoder consistently focuses on the gripper contact region, regardless of the background or whether the object is seen or unseen during training.
We further analyze how the amount of pretraining data affects reconstruction performance and representation quality. Fig. 6 shows training loss curves across different dataset sizes, demonstrating that larger datasets result in lower reconstruction losses and more stable training. To illustrate this effect, we visualize a representative reconstructed tactile image alongside its corresponding attention heatmap. As the dataset size increases, the attention maps become more sharply focused on the gripper contact regions, and the reconstructed tactile images exhibit reduced noise and clearer structure—both indicating improved cross-modal representation learning.
Figure 6: Impact of Pretraining Dataset Size on Cross-Modal Reconstruction. On the left, we show a representative visuo-tactile example with RGB image, the masked tactile input, and the ground-truth tactile image. On the right, the top panel plots test data MSE loss against the amount of pretraining demonstrations, and the bottom panel displays reconstructed tactile images alongside their visual self-attention heatmaps for four different dataset amounts. As the amount of pretraining data grows, the test MSE steadily decreases and the reconstructions become notably cleaner and more accurate.
@@ 5.2 Experimental Setup
Fluid Transfer. The robot uses a pipette to transfer water between containers. It must grasp the pipette firmly and apply just enough pressure to extract liquid without dropping it. The robot then needs to move to the top of the other container and gently squeeze to release the water into it. This task demands continuous and sensitive force modulation. Evaluation Metric: The task is considered successful if the water is transferred to the second container without spilling.
@@
(1) Vision-Only. This method feeds one RGB image as input to CLIP, and extracts a 768-dimensional CLIP embedding. This embedding, along with the robot’s proprioceptive information, is then fed into the image-based diffusion policy. We follow the same implementation as outlined in Chi et al. Chi et al. (2024).
(2) Ours w/o Cross-Attention. This method does not employ any cross-attention between vision and touch. Instead, it processes two tactile images (from the left and right fingers) through a 3-layer CNN to produce a 512-dimensional feature vector, which is then simply concatenated with the visual embedding and proprioceptive states as input to the diffusion policy.
@@ 5.3 Analysis and Discussion
(1) Tactile feedback provides explicit in-hand state information. In a single-camera handheld setup, visual inputs often suffer from severe occlusions. For example, in the Test Tube Collection task, the vision-only policy relied heavily on the color of the wooden cork to infer orientation. A minor change (such as switching to a lighter cork with less distinct features) confused the vision policy and degraded performance in reorientation. The tactile policy, however, remained unaffected by such variations.
(2) Tactile feedback improves detection of critical state transitions. In fine-grained force-controlled tasks like Fluid Transfer, accurately identifying when one action phase ends and another begins is essential. The vision-only policy struggles with this because visual features—such as the gripper’s width—remain similar before and after contact, making it difficult to determine if the gripper is still squeezing or has completed the action. As a result, the policy often skips the “expel water” phase prematurely, incorrectly assuming the transfer is finished. In contrast, the tactile policy receives direct feedback from pressure changes, enabling it to detect subtle shifts in force and correctly infer when the water has been fully expelled. The tactile feedback helps the policy transition smoothly between task stages, improving both accuracy and robustness.
(3) Joint visuo-tactile encoders enable more coordinated use of vision and touch. Naive fusion approaches that simply concatenate visual and tactile features, such as the policy without cross-attention, often fail to meaningfully combine the two modalities. As a result, the policy may over-rely on one input while neglecting the other. This imbalance was evident in the Whiteboard Erasing task: the policy without cross-attention applied excessive force to maximize tactile signal changes, which triggered safety warnings and caused the task to fail. In contrast, our jointly trained visuo-tactile encoder learns to coordinate both modalities, allowing the policy to modulate force more appropriately based on visual context and tactile feedback. This balanced integration reduces failure cases caused by overuse or misuse of either modality.
@@
To evaluate the effectiveness of pretraining, we assess the performance of the Test Tube Collection task under varying numbers of training demonstrations and epochs. The results are shown in Fig. 8. We found that pretraining provides significant benefits, particularly in low-data and low-epoch training settings.
@@
To further understand the performance gap between policies with and without pretraining observed during real-world evaluation, we analyzed the self-attention heatmaps from the ViT module in the Test Tube Collection task, as shown in Fig. 9. The policy with pretraining consistently focuses attention on the gripper–object contact region, suggesting that it retains spatial priors critical for manipulation—priors learned during large-scale visuo-tactile pretraining. In contrast, the policy without pretraining frequently directs attention to irrelevant background features, such as table edges or nearby objects, likely in an attempt to infer contact indirectly. This misplaced focus can degrade performance, particularly in cluttered or visually diverse environments. These attention patterns provide insight into why the policy with pretraining not only learns more efficiently but also generalizes better across task variations. Additional analysis of loss convergence for both policies is provided in the appendix.
@@ Acknowledgement
This work is partially supported by the Toyota Research Institute (TRI), the Sony Group Corporation, Google, Dalus AI, and an Amazon Research Award, Fall 2024. This article solely reflects the opinions and conclusions of its authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of the sponsors.
@@
Appendix A Downstream Task Training Analysis
@@
Appendix B Baseline Failure Analysis
@@
Appendix C Training Hyperparameters for Pretraining
@@
Cross-attention dropout 0.20 Weight decay $2\times 10^{-3}$
@@
Appendix D Detailed Hyperparameters for Downstream Task Training
@@
Appendix E Extended Task Descriptions
@@
E.1 Details for the Test Tube Collection Task
@@
E.2 Details for the Fluid Transfer Task
@@
E.3 Details for the Whiteboard Erasing Task
@@
E.4 Details for the Pencil Insertion Task

### v2

@@ 5.1 Large-Scale Visuo-Tactile Data and Pretraining
We evaluate the learned visuo-tactile representations through two complementary qualitative analyses. First, we test cross-modal reconstruction by providing partially masked tactile and RGB inputs and measuring the recovery of missing tactile signals, assessing whether the model captures meaningful visuo-tactile associations in both in- and out-of-distribution settings. Second, we visualize self-attention maps from the final ViT layer to examine whether the model consistently attends to contact-relevant regions in RGB images across diverse scenarios. Fig. 5 shows qualitative results from four representative tasks, including two in-distribution and two out-of-distribution examples.
Figure 5: Qualitative Results of Pretraining. We show four examples illustrating our pretrained encoder’s tactile reconstruction performance and ViT self-attention heatmaps. The encoder accurately reconstructs tactile images for both in- and out-of-distribution inputs, while the vision module consistently attends to the gripper–contact region, independent of background or object familiarity.
@@ 5.2 Experimental Setup
Fluid Transfer. The robot uses a pipette to transfer fluid between containers, requiring precise pressure to extract liquid without dropping it. It then moves above the target container and gently squeezes to release the fluid. This task demands fine-grained force control. Evaluation Metric: The task is considered successful if the fluid is transferred to the second container without spilling.
@@
(1) Vision-Only. This method feeds one RGB image as input to CLIP, and extracts a 768-dimensional CLIP embedding. This embedding, along with the robot’s proprioceptive information, is then fed into the image-based diffusion policy. We follow the same implementation as outlined in Chi et al. chi2024universal.
(2) Ours w/o Cross-Attention. This method does not employ any cross‐attention between vision and touch. Instead, it processes two tactile images (from the left and right fingers) through a 3‑layer CNN to produce a 512‑dimensional feature vector, which is then simply concatenated with the visual embedding and proprioceptive states as input to the diffusion policy.
@@ 5.3 Analysis and Discussion
(1) Tactile feedback provides explicit in-hand state information. In a single-camera setup, visual inputs often suffer from severe occlusions. For example, in the Test Tube Collection task, the vision-only policy relied heavily on the color of the wooden cork to infer orientation. A minor change (such as switching to a lighter cork with less distinct features) confused the vision policy and degraded performance in reorientation. The tactile policy, however, remained unaffected by such variations.
(2) Tactile feedback improves detection of critical state transitions. In fine-grained, force-controlled tasks like Fluid Transfer, recognizing transitions between action phases is crucial. The vision-only policy struggles because visual cues, such as gripper width, appear similar before and after contact, making it hard to tell whether squeezing has ended. Consequently, it often skips the “expel water” phase prematurely. In contrast, the tactile policy uses pressure feedback to sense subtle force changes, accurately detecting when the water is fully released. This tactile awareness enables smoother phase transitions and improves both accuracy and robustness.
(3) Joint visuo-tactile encoders enable more coordinated use of vision and touch. Naive fusion methods that simply concatenate visual and tactile features, such as the policy without cross-attention, often fail to integrate the two modalities effectively, leading the policy to over-rely on one and neglect the other. This imbalance was evident in the Whiteboard Erasing task, where the policy without cross-attention applied excessive force to amplify tactile signals, triggering safety stops. In contrast, our jointly trained visuo-tactile encoder coordinates both modalities, enabling force modulation based on visual context and tactile feedback, and reducing failures caused by misuse of either modality.
@@
To assess pretraining effectiveness, we evaluate Test Tube Collection performance across varying numbers of demonstrations and training epochs, as shown in Fig. 7. We found that pretraining provides significant benefits, particularly in low-data and low-epoch training settings.
@@ Acknowledgement
This work is partially supported by the DARPA TIAMAT program (HR0011-24-9-0430), NSF Award #2409661, Toyota Research Institute (TRI), Sony Group Corporation, Samsung Research America (SRA), Google, Dalus AI, Pickle Robot, and an Amazon Research Award (Fall 2024). This article solely reflects the opinions and conclusions of its authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of the sponsors.
@@
Contents
Appendix A Cross-Sensor Generalization and Consistency
Throughout the project, the visuo-tactile gripper and robot grippers used different pairs of tactile sensors, providing an inherent test of cross-sensor generalization. To further evaluate sensor consistency, we replaced the robot’s tactile sensors with newly fabricated ones that had not been used previously in this project and re-evaluated the Test Tube Collection task with the new sensors. As shown in Table 2, the success rates across all subtasks remained nearly identical, showing strong cross-sensor consistency and confirming that the sensors are reliable and interchangeable for downstream deployment.
Test Tube Collection Task Grasp Reorient Insert Whole Task
Policy w/ Pretraining (Original Sensors) 1.00 1.00 0.85 0.85
Policy w/ Pretraining (New Sensors) 1.00 1.00 0.85 0.85
Table 2: Policy performance with new tactile sensors. Consistent downstream performance confirms cross-sensor reliability.
Appendix B Pretraining Ablations
B.1 Effect of Demonstration Quantity on Pretraining
To assess the impact of pretraining data, we analyze how the size of the pretraining dataset influences reconstruction and representation quality in our visuo-tactile encoder. As shown in Fig. 8, larger datasets yield lower reconstruction losses and more stable training. Corresponding tactile reconstructions and ViT self-attention maps show that with more data, attention concentrates more precisely on gripper contact regions, and reconstructions become cleaner and more structured, which signals stronger cross-modal representations.
Figure 8: Impact of Pretraining Dataset Size on Cross‑Modal Reconstruction. On the left, we show a representative visuo‑tactile example with RGB image, the masked tactile input, and the ground‑truth tactile image. On the right, the top panel plots test data MSE loss against the amount of pretraining demonstrations, and the bottom panel displays reconstructed tactile images alongside their visual self‑attention heatmaps for four different dataset amounts. As the amount of pretraining data grows, the test MSE steadily decreases and the reconstructions become notably cleaner and more accurate.
B.2 Impact of Removing Task-Specific Demonstrations on Pretraining
To evaluate the robustness and generalization of our visuo-tactile pretraining, we conducted two experiments where we removed demonstrations for the specific downstream fine-tuning tasks from the pretraining dataset. In the first experiment, all demonstrations for the Test Tube Collection task were excluded, and in the second, all demonstrations for the Whiteboard Erasing task were removed. This setup allowed us to assess how effectively the pretrained visuo-tactile encoder generalizes when trained solely on in-the-wild data, without any task-specific examples included during pretraining.
After each removal, we retrained the encoder on the reduced dataset, fine-tuned a downstream policy for the corresponding task, and then evaluated its real-world performance. The resulting pretraining dataset contained approximately 2,500 in-the-wild demonstration videos, compared to the original 2,700.
Qualitatively, across both held-out tasks, we observed that the ViT attention heatmaps, evaluated on both training and held-out tasks, remained focused on the gripper’s contact regions, consistent with those obtained using the full pretraining dataset. This indicates that the encoder continued to capture meaningful visuo-tactile correspondences even when the downstream tasks were excluded from pretraining.
Quantitatively, as shown in Table 3, the downstream policies trained with and without the corresponding task data in pretraining achieved nearly identical performance. These results demonstrate that the learned visuo-tactile representations generalize effectively to unseen manipulation tasks rather than overfitting to the specific demonstrations present in the pretraining data.
Effect of Removing Task-Specific Demonstrations From Pretraining
Policy Test Tube Collection Whiteboard Erasing
Grasp Reorient Insert Whole Task First Erase Second Erase Whole Task
With Pretraining (Ours) 1.00 1.00 0.85 0.85 0.70 0.75 0.70
Without Task Data in Pretraining 1.00 1.00 0.80 0.80 0.75 0.75 0.75
Table 3: Downstream policy performance when task-specific demonstrations were excluded from pretraining. We removed all demonstrations for (1) Test Tube Collection and (2) Whiteboard Erasing from the pretraining dataset. The resulting performance remains comparable to that achieved with the full dataset, indicating that the pretrained visuo-tactile encoder generalizes well even without task-specific data.
Appendix C Downstream Task Training Analysis
C.1 Impact of Pretraining on Training Loss Convergence
@@
C.2 Policy Self-Attention Maps Before vs. After Downstream Fine-Tuning
To further understand the performance gap between policies with and without pretraining observed during real-world evaluation, we analyzed the self-attention heatmaps from the ViT module in the Test Tube Collection task, as shown in Fig. 10. The policy with pretraining consistently focuses attention on the gripper-object contact region, suggesting that it retains spatial priors critical for manipulation-priors learned during large-scale visuo-tactile pretraining. In contrast, the policy without pretraining frequently directs attention to irrelevant background features, such as table edges or nearby objects, likely in an attempt to infer contact indirectly. This misplaced focus can degrade performance, particularly in cluttered or visually diverse environments. These attention patterns provide insight into why the policy with pretraining not only learns more efficiently but also generalizes better across task variations.
Appendix D Baseline Failure Analysis
@@
Appendix E Training Hyperparameters for Pretraining
@@
Cross‐attention dropout 0.20 Weight decay $2\times 10^{-3}$
@@
Appendix F Detailed Hyperparameters for Downstream Task Training
@@
Appendix G Extended Task Descriptions
@@
G.1 Details for the Test Tube Collection Task
@@
G.2 Details for the Fluid Transfer Task
@@
G.3 Details for the Whiteboard Erasing Task
@@
G.4 Details for the Pencil Insertion Task